OpenStudy (anonymous):

During a flu epidemic in a small town, a public health official finds that the total number of people P who have caught the flu after t days is closely approximated by the formula P = −t2 + 24t + 131, where 1 ≤ t ≤ 13. (a) How many have caught the flu after 7 days? people

OpenStudy (anonymous):

replace t by 7 and compute \[-7^2+24\times 7+131\]

OpenStudy (anonymous):

i get 250 but it is late

OpenStudy (anonymous):

isn't it 549?

OpenStudy (anonymous):

oh no because don't forget that \[-7^2=-7\times 7=-49\] and not \[(-7)^2=49\]

OpenStudy (anonymous):

i checked twice, still get 250

OpenStudy (anonymous):

ohhyeah I did the second thing -_- n okey dokey thanks:D

OpenStudy (anonymous):

It's right :D:D